OpenStudy (anonymous):

(f0g) composite f(x)=3x g(x)=x+1

OpenStudy (anonymous):

f()g=f(g(x)) =3(x+1) =3x+3

OpenStudy (anonymous):

Yea! Would you mind showing me how you got that answer?

OpenStudy (anonymous):

change x by g(x) in f(x)

OpenStudy (anonymous):

Would you mind providing a visual?